OpenStudy (anonymous):

help:)) separate 114 into two parts such that when the larger is divided by the smaller , the quotient is 3 and the remainder is 18..

OpenStudy (anonymous):

a+b=114 a/b =3 with a remainder of 18. This can be written as a=3b+18 So 3b+18+b=114 4b=114-18 = 96 b=96/4=24. From a+b=114, a=90. Therefore the larger number is 90 and the smaller is 24.

OpenStudy (anonymous):